OpenStudy (anonymous):

taylor series of f(x)=sin(x), a=pi/2

OpenStudy (amistre64):

since sin(x+pi/2) = cos(x) id assume its the taylor of cos(x)

OpenStudy (anonymous):

I also need to find the radius of convergence.

OpenStudy (amistre64):

the radius is infinity

OpenStudy (anonymous):

can you tell me why/how?

OpenStudy (amistre64):

first step in constructing a series so to write out the derivative for a while; and solve them for the given a; these become your corrdinants in teh series

OpenStudy (amistre64):

essentially, your constructing a polynomial that moves the same way the function does

OpenStudy (amistre64):

when 2 functions bend and move the same (have the same derivatives) they match up along certain conditions

OpenStudy (amistre64):

\[f(x)=g(x)\text{ when } f'(x)=g'(x); ...;f^{(n)}{x}=g^{(n)}{x}\]

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

can you please help me with another if I post one more?

OpenStudy (amistre64):

i can try